# Cluster Synchronization via Graph Laplacian Eigenvectors

### v1

@@ Appendix B Single mode dynamics for weighted graphs
After this derivation, introducing weight into the system, the rest of the analysis follows [21].
Appendix C SBMs approach AEPs as size grows
@@
We consider how $E$ scales with the size of the graph in the context of the SBM model. This allows us to scale a graph, while keeping the Quotient graph associated to the QEP we consider on that graph, constant in structure (though the edge weights are subject to change). Recall that the SBM model involves a partition of the vertices such that the connection probabilities between vertices in the graph model are determined solely by the partition blocks they are assigned to. Thus choose our SBM probability matrix to be
$Pr=\begin{bmatrix}p_{11}&\cdots&p_{1r}\\ \vdots&&\vdots\\ p_{r1}&\cdots&p_{rr}\end{bmatrix}$
This means that with some choice of vertex communities, $C_{1},\dotsc,C_{r}$, we can create an adjacency matrix.
@@
Where $A^{pq}$ denotes the the adjacency block describing connections between communities $p$ and $q$. Thus every entry in this matrix has probability $p_{pq}$ of being a $1$ or a 00. The idea here is to vary these communities while keeping $Pr$ constant.
Notice that this construction allows us to create an indicator matrix $P$ for the partition that the communities provide. This allows us to define $L$ and $L^{\pi}$. Now consider $LP$. This sums the columns of $L$ whose vertices belong to the same community. This means that each block $L^{pq}$ of $L$ has its rows summed together into a column vector. Each of these entries correspond to the out degree of the vertex $i$ in community $C_{p}$ to the community $C_{q}$, which we denote as $d_{out}(i,C_{q})$. The key observation to make here is that for each vertex $i$, the connection with every vertex in $C_{q}$ is sampled independently with probability $p_{pq}$. It follows that $\mathbb{E}(d_{out}(i,C_{q}))=p_{pq}|C_{q}|$. By the law of large numbers, the values will approach this expectation with larger sample sizes. This means that in the limit of scaling $C_{q}$, the value of $d_{out}(i,C_{q})$ approaches $p_{pq}|C_{q}|$. This is true for all $i$ in $C_{p}$.
Let $d_{avg}(C_{p},C_{q})$ denote the average value of $d_{out}(i,C_{q})$ over all $i\in C_{p}$. From our previous claim, $\mathbb{E}(d_{avg}(C_{p},C_{q}))=p_{pq}|C_{q}|$ by linearity. Thus $\mathbb{E}(d_{avg}(C_{p},C_{q})-d_{out}(i,C_{q}))=0$, and consequently $\mathbb{E}(E)=\mathbb{E}(LP-PL^{\pi})=0$. Since the entries of $E$ are zero in expectation, By the law of large numbers, we can expect them to approach that value as the block sizes grow.
We mention as a note that the above result is a consequence of the structure imposed by the SBM model. Namely that its sampled graphs are equitable partitions in expectation. This is as a result of their edge sampling within blocks being of uniform probability. We also note that the above relates to the work on stochastic AEPs in [10].

### v2

@@ Appendix C Single mode dynamics for weighted graphs
After this derivation, introducing weight into the system, the rest of the analysis follows the work of KalloniatisKalloniatis (2010).
Appendix D SBMs approach AEPs in large n limit
@@
We consider how $E$ scales with the size of the graph in the context of the SBM model. This allows us to scale a graph, while keeping the Quotient graph associated to the $\delta$-QEP we consider on that graph, constant in structure (though the edge weights are subject to change). Recall that the SBM model involves a partition of the vertices such that the connection probabilities between vertices in the graph model are determined solely by the partition blocks they are assigned to. Thus choose our SBM probability matrix to be
$Pr=\begin{bmatrix}\rho_{11}&\cdots&\rho_{1r}\\ \vdots&&\vdots\\ \rho_{r1}&\cdots&\rho_{rr}\end{bmatrix}$
This means that with some choice of vertex communities, $C_{1},\dotsc,C_{r}$, we can sample an adjacency matrix.
@@
Where $A^{pq}$ denotes the the adjacency block describing connections between communities $p$ and $q$. Every entry in $A^{pq}$ with $p\neq q$ has probability $\rho_{pq}$ of being a $1$ and is otherwise zero. The diagonal blocks $A^{pp}$ have the additional requirements of being symmetric and maintaining zeros on their diagonal. We vary these communities while keeping $Pr$ constant.
This construction allows us to create an indicator matrix $P$ for the partition that the communities provide. This allows us to define $L$ and $L^{\pi}$. Now consider $LP$. This sums the columns of $L$ whose vertices belong to the same community. This means that each block $L^{pq}$ of $L$ has its rows summed together. Each of these entries correspond to the out degree of the vertex $i$ in community $C_{p}$ to the community $C_{q}$, which we denote as $d_{out}(i,C_{q})$. The key observation to make here is that for each vertex $i$, the connection with every vertex in $C_{q}$ is sampled independently with probability $\rho_{pq}$. It follows that $\mathbb{E}(d_{out}(i,C_{q}))=\rho_{pq}|C_{q}|$. By the law of large numbers, the values will approach this expectation with larger sample sizes. This means that in the limit of scaling $C_{q}$, the value of $d_{out}(i,C_{q})$ approaches $\rho_{pq}|C_{q}|$. This is true for all $i$ in $C_{p}$.
Let $d_{avg}(C_{p},C_{q})$ denote the average value of $d_{out}(i,C_{q})$ over all $i\in C_{p}$. From our previous claim, $\mathbb{E}(d_{avg}(C_{p},C_{q}))=\rho_{pq}|C_{q}|$ by linearity. Thus $\mathbb{E}(d_{avg}(C_{p},C_{q})-d_{out}(i,C_{q}))=0$, and consequently $\mathbb{E}(E)=\mathbb{E}(LP-PL^{\pi})=0$. Since the entries of $E$ are zero in expectation, By the law of large numbers, we can expect them to approach that value as the block sizes grow.
We mention as a note that the above result is a consequence of the structure imposed by the SBM model. Namely that its sampled graphs are equitable partitions in expectation. This is as a result of their edge sampling within blocks being of uniform probability. We also note that the above relates to the work of Schaub et al. on stochastic AEPs.Schaub, Li, and Peel (2023)